# Learning entanglement from tomography data: contradictory measurement importance for neural networks and random forests

### v1

@@ I Introduction
For mixed states, even theoretical determination of entanglement becomes challenging and it’s experimental studies require quantum state tomography (QST) of the whole density matrix [1] followed by the calculation of one of the more numerically accessible entanglement measures. Since quantum state tomography requires $\mathcal{N}^{2}-1$ measurements, where $\mathcal{N}$ is the dimension of the whole system, the number of required measurements grows rapidly with system size (i.e. $\mathcal{N}=2^{\mathrm{\#\,qubits}}$) and becomes unmanageable for pretty small systems. For larger systems quantum state reconstruction can be effectively (with learned positivity constraint) performed using deep neural networks (NNs) [2, 3, 4, 5, 6, 7, 8], that can converge much faster [7] than the standard QST.
@@ V.2 Deep neural networks
The second type of ML model employed is the NN MLP model. MLPs are supervised models that have been successfully applied to both classification and regression tasks. In this work, we employed an MLP to accurately predict the value of concurrence $C$ using measurement values as inputs, as presented in Fig. 2(a). Our MLP architecture comprises just two fully-connected hidden layers, each containing $128$ units (neurons), and utilizing a ReLU activation function [38]. Finally, it has a single output unit with a linear activation function that predicts the $C$ value. We limited the model to only 2 hidden layers, which proved sufficient for making 2-qubit predictions. However, we also trained a much larger model consisting of 10 hidden layers, but it achieved comparable results.
The MLP is trained using the training dataset, ${\bm{S}}_{\mathrm{train}}$, (of size $N$) using a standard mean square error (MSE) loss function,

### v2

@@ I Introduction
For mixed states, even theoretical determination of entanglement becomes challenging and it’s experimental studies require quantum state tomography (QST) of the whole density matrix [1] followed by the calculation of one of the more numerically accessible entanglement measures. Since quantum state tomography requires $\mathcal{N}^{2}-1$ measurements, where $\mathcal{N}$ is the dimension of the whole system, the number of required measurements grows rapidly with system size (i.e. $\mathcal{N}=2^{\mathrm{\#\,qubits}}$) and becomes unmanageable for pretty small systems. For larger systems quantum state reconstruction can be effectively (with learned positivity constraint) performed using deep neural networks (NNs) [2, 3, 4, 5, 6, 7, 8, 9], that can converge much faster [7] than the standard QST, even for noisy measurement data [9].
@@ V.2 Deep neural networks
The second type of ML model employed is the NN MLP model. MLPs are supervised models that have been successfully applied to both classification and regression tasks. In this work, we employed an MLP to accurately predict the value of concurrence $C$ using measurement values as inputs, as presented in Fig. 2(a). Our MLP architecture comprises just two fully-connected hidden layers, each containing $128$ units (neurons), and utilizing a ReLU activation function [39]. Finally, it has a single output unit with a linear activation function that predicts the $C$ value. We used a linear output activation, although a sigmoid function may seem more natural. However, combining mean square error (MSE) as the loss function with a sigmoid activation often leads to optimization issues, as the resulting cost surface becomes non-convex. To ensure that the predictions for $C$ remained within its natural range, we applied additional clipping. We limited the model to only 2 hidden layers, which proved sufficient for making 2-qubit predictions. However, we also trained a much larger model consisting of 10 hidden layers, but it achieved comparable results.
The MLP is trained using the training dataset, ${\bm{S}}_{\mathrm{train}}$, (of size $N$) using a standard MSE loss function,
@@ VI.2.2 Measurement importance via perturbations
In addition, in Appendix E we compare the average performance of RF and NN models under a single measurement perturbation. It is shown, that NN model dominates at lower noise levels. However, as the noise amplitude increases, the the accuracy of NN model decreases faster than the one of RF. As a result, the RF model outperforms the NN at higher noise level (see Fig. 10). RF models combine the predictions of many decision trees trained on bootstrap‑resampled subsets of data. Because each tree sees a different slice of the input features, their averaged vote smooths out individual errors, making the overall model noticeably sturdier against noisy features and labels.
@@ VII Conclusions
On the other hand, the RF models can outperform the networks in the presence of noise perturbing a single measurement, as demonstrated in Appendix E. The higher robustness of RF model to the noise perturbation stems from the cooperation of a big number of classifiers relying on different features of the input vector.
@@ Appendix D Shapley interaction values for RFs
Appendix E Perturbation measure
Figure 10: Perturbation measure for (a) random forest classifier, and (b) neural network as a function of the noise amplitude $\sigma$.
In addition, we compare the performance of random forest and neural network for random noise applied to a single measurement outcome, as described in section VI.2.2. Fig. 10 shows accuracy calculated for RF1 and NN models. Although in case of zero noise, the accuracy of NN exceeds the one of RF model, in the presence of certain noise level the RF model might in average perform better.
The reason for this behavior is the large number of classifiers in the RF ensemble, which makes the method more robust against noise and errors. Because each tree in the RF ensemble classifies a sample in a different way, a random change of one feature is unlikely to affect prediction of every tree in the ensemble. Therefore, even stronger noise in the data does not significantly affect the overall performance of the model. On the other hand, NN is a highly nonlinear model, which often relies not just on the single parameters, but also on their correlations. Thus, a systematic perturbation of one of the correlated features might significantly reduce the model accuracy.